# Diversity of rocky planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v1

@@ 4.1 Earth-sized planets
Condensates at planetary surfaces, such as water, graphite, -sulfur, and ammonium chloride, are of primary interest for the establishment of geochemical cycles (Hirschmann, 2018; Dasgupta et al., 2024) and a habitable environment (Benner et al., 2020; Mrnjavac et al., 2023; Rimmer & Shorttle, 2024). Figures 9 and A5 show atmospheres at 280 K above surfaces of an Earth-like planet together with stable condensates with which they are in equilibrium, derived from a fully molten and partially molten mantle, respectively. Each panel mandates that a particular condensate is stable with a minimum mass (shown on the x-axis), where the upper row reveals the number of models associated with a different dominant (VMR50%) species as a function of the number of Earth-ocean equivalents (EOs) condensed. The lower row shows the variety of atmospheric species for dominant classes of atmospheres that support condensates, where either CO2 or CH4 is chosen as the dominant species (Table 3). The conditions for water oceans and graphite to be stable are comparable, with two classes of atmospheres emerging as determined by a volume mixing ratio greater than 50%: CO2-rich and CH4-rich, with the former representing 83% of all atmospheres, whereas the latter only around 5% (e.g., Figure 9). For both classes, the atmospheres have median total surface pressures of a few tens of bar, and due to buffering imposed by C(cr) and H2O(l) condensates, the partial pressures of other C- and H-bearing species are relatively constant, regardless of the condensate reservoir mass.
@@
Small amounts of water can be stable beneath atmospheres dominated by SO2, N2, H2, and ClH, although these are all less prevalent than CO2 and CH4 (Figure 9a). Conditions that permit the condensation of solid ammonium chloride, ClH4N(cr) mirror those for water and graphite, reflecting the greater availability of NH3 and HCl in reduced atmospheres relative to their oxidised counterparts (Figure 2). Ammonium chloride most commonly condenses in small mass fractions (0.0025 EOs) beneath CO2-rich atmospheres, but the highest ClH4N(cr) mass fractions (0.006 up to 0.012 EOs) occur, albeit rarely (1% of all models), in equilibrium with H2. Conversely, -sulfur stability occurs almost exclusively beneath CO2-rich atmospheres, in which the partial pressure of CO2 is on average higher than in equilibrium with water and graphite (compare orange line in Figure 9e and g). The CO2-rich atmospheres in equilibrium with -sulfur also frequently have several tens of bar of SO2 and only 10-11 bar of H2S. Note that SO2(g) is almost never the major gas phase above -sulfur, as most of the S budget is hosted in the condensed phase. In total, around 40% of models have co-existence of water, graphite, -sulfur, and ammonium chloride, albeit at potentially low abundance. Nevertheless, our results show that many early atmospheres are conducive to the formation of condensates at the planetary surface, should the planetary surface reach its equilibrium temperature.
@@
Figure 10 summarises the atmospheric abundances (ppm) above planetary surfaces that simultaneously support water, graphite, -sulfur, and ammonium chloride. The objective is to identify characteristic atmospheric signatures that may be used to identify planetary surfaces with a higher propensity for habitability (Krissansen-Totton et al., 2018; Schwieterman et al., 2019; Konrad et al., 2024; Tokadjian et al., 2024; Borges et al., 2024; Young et al., 2024). To satisfy joint condensate stability, reduced species, such as H2, NH3 and CH4 are low in abundance—almost always below the ppm level—because otherwise -sulfur does not form. This comes from the fact that these species are present in the atmosphere only at very low O2 at the high temperature initial conditions, at which S is highly soluble. Therefore, S is never present in sufficient amounts in these atmospheres to be able to condense upon cooling to a low temperature. Conversely, oxidised initial conditions that produce significant SO2 lead to final, elevated O2s of these atmospheres that prevent the condensation of graphite via CO2(g) = C(cr) + O2(g), even at high CO2. Consequently, any detection of SO2 in an atmosphere is incompatible with the presence of graphite, and is highly unlikely to be the dominant atmospheric gas in the presence of H2O(l) or -S(cr). Species whose fugacities are less redox sensitive cannot readily be used to discriminate between stable versus unstable condensates (e.g., H2O and N2). Figure A6 illustrates the same analysis but excludes -sulfur and ammonium chloride and hence only considers the joint stability of water and graphite. In this scenario, reduced species can be accommodated at higher abundances (e.g. H2, NH3), and notably CH4-rich atmospheres are compatible with both water oceans and graphite at planetary surfaces.
@@
The establishment of geological cycling may require an early reservoir of water at the surface or in the upper mantle (e.g., Tikoo & Elkins-Tanton, 2017) to alter the rheology and partial melting of rocks (Korenaga, 2020). In this regard, surfaces that emerge post-formation with condensates, notably liquid water, may already be primed to cycle volatiles with the interior and thereby modify the size and composition of the atmosphere. Conversely, for surfaces that emerge without condensates, there could be a higher propensity to remain locked in an inactive state since geochemical cycling cannot initiate, at least not immediately. Table 3 reveals that most Earth-like planets that emerge from the magma ocean stage have the prerequisite conditions for surficial water, with at least 56% of planets also accommodating graphite. Since condensates buffer the atmospheric speciation, finding atmospheres that are compatible with stable condensates will not directly allow the size of those condensate reservoirs to be determined.
@@ 4.2.1 Non-ideality and the ‘fugacity crisis’
Regarding point 1, Kite et al. (2019) proposed that a ‘fugacity crisis’, in which the enhanced dissolution of H2 at high surface pressures (1–8 GPa) hinders atmospheric growth, can explain the drop-off in exoplanet abundance above 3 R⊕ observed in exoplanet demographics studies (i.e., the ‘radius cliff’) (Fulton & Petigura, 2018); however, see also Schlichting & Young (2022) for a different interpretation. Our work goes beyond the H2 non-ideality case considered in Kite et al. (2019), by incorporating additional volatile species (e.g., H2O, CH4, CO and CO2) along with their dissolution into the interior and real gas EOS. Despite the relatively large mass of envelopes (2–3% of the planetary mass) required to explain the mass-radius relationships of sub-Neptune exoplanets in the absence of interior dissolution, these quantities are modest, on a molar basis, with respect to the proportion of condensed mass in the planet. As such, in the end-member case of a solar gas interacting with a rocky sub-Neptune interior, the equilibrium H/O and C/O ratios will be modified from their solar values (2000 and 0.55 by moles, respectively, Asplund et al., 2009) via the mutual solubility of O in the envelope and C and H in the mantle. Since the H/O and C/O ratios of nominally rocky material are 1 (H/O and C/O ratios in the bulk silicate Earth are 0.005 and 3.6 10-4 by moles, respectively, Palme & O’Neill, 2014), chemical equilibrium will result in a decrease in H/O and C/O in the envelope relative to solar values (see also Young et al., 2023).
Here, because only volatile element mass balance is considered (that is, there is no notion of the chemical composition of the condensed phase other than for C, H, O and N), this process is simulated by considering a higher O2 (IW=-3) in our sub-Neptune models than that imposed by the solar nebula gas (IW=-6.5, Grossman et al., 2008). Should O2 at the magma ocean-envelope interface fall below IW=-3.5, then H2 predominates in the gas phase (envelope), even for C/H ratios 100 solar (Figure F1). As noted by Kite et al. (2019); Schlichting & Young (2022), however, the experiments underpinning the solubility law for H2 dissolution in silicate liquids were performed over a limited temperature (1673–1773 K) and pressure (up to 3 GPa) range (Hirschmann et al., 2012); the temperature range is notably much lower than the 3000 K assumed for equilibrium at the interface. Therefore, how the equilibrium constant of the reaction H2(g) = H2(l) depends on temperature, and pressures above 3 GPa, is unconstrained, yet is essential to assess the ‘fugacity crisis’ hypothesis.
@@ 4.2.2 Role of non-ideality in inferring atmospheric chemistry
A simplified version of this exercise was carried out by Yang & Hu (2024), in which speciation at the pressures of observation in the atmospheres of K2-18b and TOI-270d were extrapolated to 200 bar. In order to match the CH4/CO2 proportions in the upper atmosphere, these authors inferred H2/H2O ratios at 200 bar to be roughly unity for K2-18b and 0.5 for TOI-270d. Assuming these ratios are representative of those at any magma ocean/atmosphere interface, then this would correspond to IW in the range 0 to 1. These O2s are similar to, or more oxidised than, those inferred for small, differentiated rocky bodies in the solar system (Wadhwa, 2008). Such high oxygen fugacities would seem at odds with the substantial quantities (1–3% by mass) of solar-like gas inferred to have been accreted by sub-Neptunes. Indeed, we predict C/H ratios that are too high (2.5) at these O2s, and H2 would no longer be a predominant constituent of the atmosphere (Figure F1). Therefore, preliminary indications suggest that either 1) bulk planetary C/H ratios should be relatively low (10-1 by mass), either by processes that fractionate C from other atmosphere-forming elements during accretion relative to solar or its preferential sequestration into sub-Neptune cores and/or 2) the lower atmosphere has a distinct chemical composition with higher C/H and C/O than does the upper atmosphere sensed by transmission measurements (e.g., by inefficient mixing or condensation of graphite in the atmospheric column), in order to maintain simultaneously low C/H (0.1–0.3) and C/O (0.02–0.5) ratios observed in temperate sub-Neptunes.
@@ 5 Conclusions
For Earth-sized rocky planets like TRAPPIST-1e, the atmosphere during the magma ocean stage is likely CO-rich due to two key factors. First, CO is thermodynamically favoured at 1800 K across a broad range of oxygen fugacities. Second, carbon-bearing species are significantly less soluble than hydrogen-bearing species—particularly H2O—leading to an elevated atmospheric C/H, often exceeding the bulk C/H by several orders of magnitude. As a result, hydrogen-bearing species are depleted in the atmosphere, even when the mantle is only partially molten. In a fully molten mantle, the further depletion of hydrogen sequesters H2O and leads to atmospheres that are defined by entirely continuous mixing ratios of their constituent species, favouring atmospheres dominated by 1) heavy carbon species (predominantly CO), 2) CO2 SO2-bearing varieties at high O2s (IW=+3 to +4) or 3) H2 CH4-bearing types at low O2 (below IW=-3).
Based on the equilibrium results of atmospheres with molten mantles, we cool the atmospheres to 280 K and permit (if thermodynamically stable) the formation of water oceans, graphite, -sulfur, and ammonium chloride. Unlike at high temperatures, the presence of condensates buffers gas mixing ratios to produce discrete ‘families’ of atmospheres. Notably, CO2-rich atmospheres are not only widespread (83% of cases) but also uniquely capable of supporting a wider variety of surface condensates compared to those dominated by other species. A small fraction (5%) are CH4 atmospheres that can sustain both water and graphite while roughly 8% are CO2-SO2-rich in equilibrium with native sulfur. The remainder (1%) are H2-rich. Depending on the conditions during early equilibration with a magma ocean, atmospheres rich in N2, ClH, or H2S can also host diverse surface condensates; however, such cases compromise only 1% of all cases. In short, water condensation buffers the partial pressures of hydrogen-bearing species in the atmosphere, such that even low C/H does not produce substantial hydrogen atmospheres except at the most reduced conditions where H2 can dominate. Thus, for both high and low temperature atmospheres, carbon always wins.
@@ Acknowledgements
Appendix A Partially molten mantle

### v2

@@ 4.1 Earth-sized planets
Condensates at planetary surfaces, such as water, graphite, -sulfur, and ammonium chloride, are of primary interest for the establishment of geochemical cycles (Hirschmann, 2018; Dasgupta et al., 2024) and a habitable environment (Benner et al., 2020; Mrnjavac et al., 2023; Rimmer & Shorttle, 2024). Figures 9 and B5 show atmospheres at 280 K above surfaces of an Earth-like planet together with stable condensates with which they are in equilibrium, derived from a fully molten and partially molten mantle, respectively. Each panel mandates that a particular condensate is stable with a minimum mass (shown on the x-axis), where the upper row reveals the number of models associated with a different dominant (VMR50%) species as a function of the number of Earth-ocean equivalents (EOs) condensed. The lower row shows the variety of atmospheric species for dominant classes of atmospheres that support condensates, where either CO2 or CH4 is chosen as the dominant species (Table 3). The conditions for water oceans and graphite to be stable are comparable, with two classes of atmospheres emerging as determined by a volume mixing ratio greater than 50%: CO2-rich and CH4-rich, with the former representing 83% of all atmospheres, whereas the latter only around 5% (e.g., Figure 9). For both classes, the atmospheres have median total surface pressures of a few tens of bar, and due to buffering imposed by C(cr) and H2O(l) condensates, the partial pressures of other C- and H-bearing species are relatively constant, regardless of the condensate reservoir mass.
@@
Small amounts of water can be stable beneath atmospheres dominated by SO2, N2, H2, and ClH, although these are all less prevalent than CO2 and CH4 (Figure 9a). Conditions that permit the condensation of solid ammonium chloride, ClH4N(cr) mirror those for water and graphite, reflecting the greater availability of NH3 and HCl in reduced atmospheres relative to their oxidised counterparts (Figure 2). Ammonium chloride most commonly condenses in small mass fractions (0.0025 EOs) beneath CO2-rich atmospheres, but the highest ClH4N(cr) mass fractions (0.006 up to 0.012 EOs) occur, albeit rarely (1% of all models), in equilibrium with H2. Conversely, -sulfur stability occurs almost exclusively beneath CO2-rich atmospheres, in which the partial pressure of CO2 is on average higher than in equilibrium with water and graphite (compare orange line in Figure 9e and g). The CO2-rich atmospheres in equilibrium with -sulfur also frequently have several tens of bar of SO2 and only 10-11 bar of H2S. Note that SO2(g) is almost never the major gas phase above -sulfur, as most of the S budget is hosted in the condensed phase. In total, water, graphite, -sulfur, and ammonium chloride coexist in around 40% of the models, albeit at potentially low abundance. Nevertheless, our results show that many early atmospheres are conducive to the formation of condensates on temperate planetary surfaces (here 280 K), should these atmospheres be able to cool isochemically and reach chemical equilibrium with stable condensates.
@@
Figure 10 summarises the atmospheric abundances (ppm) above planetary surfaces that simultaneously support water, graphite, -sulfur, and ammonium chloride. The objective is to identify characteristic atmospheric signatures that may be used to identify planetary surfaces with a higher propensity for habitability (Krissansen-Totton et al., 2018; Schwieterman et al., 2019; Konrad et al., 2024; Tokadjian et al., 2024; Borges et al., 2024; Young et al., 2024a). To satisfy joint condensate stability, reduced species, such as H2, NH3 and CH4 are low in abundance—almost always below the ppm level—because otherwise -sulfur does not form. This comes from the fact that these species are present in the atmosphere only at very low O2 at the high temperature initial conditions, at which S is highly soluble. Therefore, S is never present in sufficient amounts in these atmospheres to be able to condense upon cooling to a low temperature. Conversely, oxidised initial conditions that produce significant SO2 lead to final, elevated O2s of these atmospheres that prevent the condensation of graphite via CO2(g) = C(cr) + O2(g), even at high CO2. Consequently, any detection of SO2 in an atmosphere is incompatible with the presence of graphite. On the other hand, SO2 is highly unlikely to be the dominant atmospheric gas in the presence of H2O(l) or -S(cr). Species whose fugacities are less redox sensitive cannot readily be used to discriminate between stable versus unstable condensates (e.g., H2O and N2). Figure B6 illustrates the same analysis but excludes -sulfur and ammonium chloride and hence only considers the joint stability of water and graphite. In this scenario, reduced species can be accommodated at higher abundances (e.g. H2, NH3), and notably CH4-rich atmospheres are compatible with both water oceans and graphite at planetary surfaces.
@@
The establishment of geological cycling may require an early reservoir of water at the surface or in the upper mantle (e.g., Tikoo & Elkins-Tanton, 2017) to alter the rheology and partial melting of rocks (Korenaga, 2020). In this regard, surfaces that emerge post-formation with condensates, notably liquid water, may already be primed to cycle volatiles with the interior and thereby modify the size and composition of the atmosphere. Conversely, for surfaces that emerge without condensates, there could be a higher propensity to remain locked in an inactive state since geochemical cycling cannot initiate, at least not immediately. Table 3 reveals that most (91%) Earth-like planets that emerge from the magma ocean stage have the prerequisite conditions for surficial water, with at least 56% of planets also accommodating graphite. Since condensates buffer the atmospheric speciation (due to the equivalence of chemical potentials, an intensive variable), finding atmospheres that are compatible with stable condensates will not directly allow the size of those condensate reservoirs to be determined.
@@ 4.2.1 Non-ideality and the ‘fugacity crisis’
Regarding point 1, Kite et al. (2019) proposed that a ‘fugacity crisis’, in which the enhanced dissolution of H2 at high surface pressures (1–8 GPa) hinders atmospheric growth, can explain the drop-off in exoplanet abundance above 3 R⊕ observed in exoplanet demographics studies (i.e., the ‘radius cliff’) (Fulton & Petigura, 2018); though see also Schlichting & Young (2022) for a different interpretation. Our work goes beyond the H2 non-ideality case considered in Kite et al. (2019), by incorporating additional volatile species (e.g., H2O, CH4, CO and CO2) along with their dissolution into the interior and real gas EOS. Despite the relatively large mass of envelopes (2–3% of the planetary mass) required to explain the mass-radius relationships of sub-Neptune exoplanets in the absence of interior dissolution, these quantities are modest, on a molar basis, with respect to the proportion of condensed mass in the planet. As such, in the end-member case of a solar gas interacting with a rocky sub-Neptune interior, the equilibrium H/O and C/O ratios will be modified from their solar values (2000 and 0.55 by moles, respectively, Asplund et al., 2009) via the mutual solubility of O in the envelope and C and H in the mantle. Since the H/O and C/O ratios of nominally rocky material are 1 (H/O and C/O ratios in the bulk silicate Earth are 0.005 and 3.6 10-4 by moles, respectively, Palme & O’Neill, 2014), chemical equilibrium will result in a decrease in H/O and C/O in the envelope relative to solar values (see also Young et al., 2023).
Here, because only volatile element mass balance is considered (that is, there is no notion of the chemical composition of the condensed phase other than for C, H, O and N), this process is simulated by considering a higher O2 (IW=-3) in our sub-Neptune models than that imposed by the solar nebula gas (IW=-6.5, Grossman et al., 2008). Should O2 at the magma ocean-envelope interface fall below IW=-3.5, then H2 predominates in the gas phase (envelope), even for C/H ratios 100 solar, though CH4 remains a minor species (Figure G1). As noted by Kite et al. (2019); Schlichting & Young (2022), however, the experiments underpinning the solubility law for H2 dissolution in silicate liquids were performed over a limited temperature (1673–1773 K) and pressure (up to 3 GPa) range (Hirschmann et al., 2012); the temperature range is notably much lower than the 3000 K assumed for equilibrium at the interface. Therefore, how the equilibrium constant of the reaction H2(g) = H2(l) depends on temperature, and pressures above 3 GPa, is unconstrained, yet is essential to assess the ‘fugacity crisis’ hypothesis.
@@ 4.2.2 Role of non-ideality in inferring atmospheric chemistry
A simplified version of this exercise was carried out by Yang & Hu (2024), in which speciation at the pressures of observation in the atmospheres of K2-18b and TOI-270d were extrapolated to 200 bar. In order to match the CH4/CO2 proportions in the upper atmosphere, these authors inferred H2/H2O ratios at 200 bar to be roughly unity for K2-18b and 0.5 for TOI-270d. Assuming these ratios are representative of those at any magma ocean/atmosphere interface, then this would correspond to IW in the range 0 to 1. These O2s are similar to, or more oxidised than, those inferred for small, differentiated rocky bodies in the solar system (Wadhwa, 2008). Such high oxygen fugacities would seem at odds with the substantial quantities (1–3% by mass) of solar-like gas inferred to have been accreted by sub-Neptunes. Indeed, we predict C/H ratios that are too high (2.5) at these O2s, and H2 would no longer be a predominant constituent of the atmosphere (Figure G1). Therefore, preliminary indications suggest that either 1) bulk planetary C/H ratios should be relatively low (10-1 by mass), either by processes that fractionate C from other atmosphere-forming elements during accretion relative to solar or its preferential sequestration into sub-Neptune cores and/or 2) the lower atmosphere has a distinct chemical composition with higher C/H and C/O than does the upper atmosphere sensed by transmission measurements (e.g., by inefficient mixing or condensation of graphite in the atmospheric column), in order to maintain simultaneously low C/H (0.1–0.3) and C/O (0.02–0.5) ratios observed in temperate sub-Neptunes.
@@ 5 Conclusions
For Earth-sized rocky planets like TRAPPIST-1e, the atmosphere during the magma ocean stage is likely CO-rich due to two key factors. First, CO(g) is thermodynamically favoured at 1800 K across a broad range of oxygen fugacities. Second, carbon-bearing species are significantly less soluble than hydrogen-bearing species—particularly H2O(g)—leading to an elevated atmospheric C/H, often exceeding the bulk C/H by several orders of magnitude. As a result, hydrogen-bearing species are depleted in the atmosphere, even when the mantle is only partially molten. In a fully molten mantle, the further depletion of hydrogen by the sequestration of H2O(g) into the liquid leads to atmospheres that are defined by entirely continuous mixing ratios of their constituent species, favouring atmospheres dominated by 1) heavy carbon species (predominantly CO(g)), 2) CO2(g) SO2(g)-bearing varieties at high O2s (IW=+3 to +4) or 3) H2(g) CH4(g)-bearing types at low O2 (below IW=-3).
Based on the equilibrium results of atmospheres with molten mantles, we cool the atmospheres to 280 K and permit (if thermodynamically stable) the formation of water (oceans), graphite, -sulfur, and ammonium chloride. Unlike at high temperatures, the presence of condensates buffers gas mixing ratios to produce discrete ‘families’ of atmospheres. Notably, CO2-rich atmospheres are not only widespread (83% of cases) but also uniquely capable of supporting a wider variety of surface condensates compared to those dominated by other species. A small fraction (5%) are CH4-rich atmospheres that can sustain both water and graphite while roughly 8% are CO2-SO2-rich in equilibrium with native sulfur. The remainder (1%) are H2-rich. Depending on the conditions during early equilibration with a magma ocean, atmospheres rich in N2, HCl, or H2S can also host diverse surface condensates; however, such cases comprise only 1% of all cases. In short, water condensation buffers the partial pressures of hydrogen-bearing species in the atmosphere, such that even low C/H does not produce substantial hydrogen atmospheres except at the most reduced conditions where H2 can dominate. Thus, for both high and low temperature atmospheres, carbon always wins.
@@ Acknowledgements
Appendix A Sampling strategy and parameter choices
The logarithmic sampling strategy adopted to explore the atmospheric diversity of TRAPPIST-1e is physically motivated as follows. Oxygen fugacity (O2) exerts a fundamental control on atmospheric composition, particularly through its influence on redox speciation, for example, determining the relative abundances of reduced versus oxidised species such as H2 and H2O. Thermodynamically, O2 is defined by an equilibrium reaction, and when referenced to Fe–FeO (iron-wüstite, IW) equilibrium (Equation E1) is conventionally expressed as a offset relative to the IW buffer. This convention naturally supports the use of a log-uniform prior, especially within a thermochemical framework where partial pressures, fugacities, and equilibrium constants combine multiplicatively (and thus additively in log-space). Observationally, the present-day O2s of rocky bodies in the solar system spans several orders of magnitude from roughly 5 units below IW (Mercury) to 2 units above IW (carbonaceous chondrites) (e.g., Grewal et al., 2024), with several extrasolar rocky bodies showing comparable redox states (Doyle et al., 2019). Early Earth conditions during the magma ocean stage were close to IW (Sossi et al., 2020).
Therefore, to capture this diversity, we centre our log-uniform prior on zero offset relative to IW (anchored to early Earth) and allow sampling across a broad range of units. While no prior is without limitations, we adopt this scale-invariant formulation because it more faithfully represents ignorance over the parameter’s order of magnitude. It offers the dual advantages of anchoring to a well-motivated reference point (early Earth) while permitting exploration beyond the present-day solar system constraints on O2, which is warranted by the potentially greater diversity of exoplanetary environments. For consistency with the log-uniform prior, we apply the same sampling strategy to the other two parameters: the total hydrogen inventory and the C/H mass ratio are each varied across one order of magnitude on either side of an Earth-like reference value. The upper bound of the total hydrogen inventory is motivated by planet formation simulations, which show that terrestrial planets can accrete several times Earth’s water content (Raymond et al., 2007). The lower bound, by contrast, reflects scenarios of limited delivery or substantial early loss.
The corner plots display bivariate scatter plots (e.g., Figure 1), which directly reflect the model outcomes and are therefore insensitive to the choice of prior, depending only on the sampled parameter ranges. In contrast, the likelihood contours, marginal distributions, and derived summary statistics (e.g., Table 3) are prior-dependent. A uniform distribution might appear to be a simple alternative prior, but its inadequacy for O2 is evident: when sampled uniformly between and , the mean value in space is , heavily skewed toward the upper bound and far outside the range characteristic of the early Earth or any known rocky body in the solar system. New observations and models may eventually reveal the true distributions of the parameters that govern atmospheric diversity, but in the absence of such knowledge, a logarithmic sampling strategy remains the most reasonable and defensible choice.
Appendix B Partially molten mantle